# How to Learn a Star: Binary Classification with Starshaped Polyhedral Sets

### v1

@@ 3.1 Geometry of the 0/1-loss
We seek to understand the geometry inside the parameter space $\mathbb{R}_{>0}^{n}=\{\mathbf{a}:\mathbf{a}>0\}$. For an example which illustrates all definitions and results stated in this and the following subsection (Sections 3.1 and 3.2), we refer to Section A.2 in Section A.2.
@@
A direct consequence of Section 3.1 is that the false positives, the false negatives and thus also the discrete loss is constant on the half-open chambers.
In general, since every chamber is convex and corresponds to a unique classification, the set of all parameters $\mathbf{a}$ that perfectly separate the data points is convex.
[…]
@@
Section 3.1 implies that if the dataset is separable, i.e., if $L(\operatorname{err},0)\neq\emptyset$, then the set of perfect classifiers $L(\operatorname{err},0)$ is a convex set. Under the same assumption, we can make a similar statement as Theorem 3.2 about the sublevel sets of the discrete loss function.
@@
Theorem 3.3 shows that if the data is separable, then the sublevel sets of $\operatorname{err}$ are star-convex, but not necessarily convex in the usual sense (see also Section A.2 in Section A.2). Much stronger, in the case of non-separable data, sublevel sets and even the set of minimizers of the discrete loss can be disconnected. One example for this scenario is the point configuration from Section A.2, with labeling $0,0,1,1,1,1,0,0$. Here, the minimum of the discrete loss-function is $4$, attained in four non-neighboring half-open chambers.
3.2 Exponential loss
In Section 3.1 we have observed that the 0/1-loss or discrete loss admits discrete geometric structures in the parameter space which are governed by an affine hyperplane arrangement. However, the discrete loss function is difficult to compute in practice. We thus propose an alternative loss-function for practical purposes. The exponential loss function turns out to be well-suited for optimization procedures due to concavity.
@@ 3.2 Exponential loss
The definitions and results stated in this subsection are exemplified in Section A.2 in Section A.2. The computation of the maximum likelihood estimator, the maximizer of the exponential loss-function, can be summarized by the following algorithm.
@@
Convex (and thus also concave) functions can be optimized in polynomial time. Moreover, the description in Section 2 implies that $A_{X}$ can be determined in polynomial time. This implies the following.
[…]
Algorithm 1 can be computed in polynomial time in the size of the input data.
Similarly to Section 3.1, we now consider the superlevel sets of the exponential loss function. For given $t\in\mathbb{R}$, the superlevel $S_{\mathcal{L}}(t)$ of $\mathcal{L}$ is defined as
@@
We now analyze cases in which the maximum of the log-likelihood function is unique. For this, consider the positively labeled subdataset $X_{1}=\{\mathbf{x}^{(i)}\colon y^{(i)}=1\}\subset X$ and let $A_{X_{1}}$ be the submatrix of $A_{X}$ composed of all rows $[\mathbf{x}^{(i)}]$ corresponding to $\mathbf{x}^{(i)}\in X_{1}$; similarly we define $A_{X_{0}}$ for $X_{0}=X\setminus X_{1}$. Then we have the following sufficient condition for a unique maximum of $\mathcal{L}(\mathbf{a})$.
@@
The exponential loss function $\mathcal{L}(\mathbf{a})$ is strictly concave if the matrix $A_{X_{1}}$ has rank $n$. Furthermore, if also the rank of $A_{X_{0}}$ equals $n$, then $\mathcal{L}(\mathbf{a})$ has a unique (possibly degenerate) maximum in the closed positive orthant $\mathbb{R}_{\geq 0}^{n}$.
Matrices of the form $A_{X}$ also appear in the study of reconstruction of polytopes with fixed facet directions from support function evaluations. Dostert and Jochemko (2023) showed that such a reconstruction of a polytope is unique if and only if $\operatorname{rank}A_{X}=n$. In other words, $A_{X}\in(\mathbb{R}^{n\times m})\setminus\mathbb{V}$ where $\mathbb{V}$ is the algebraic variety encoding that $\operatorname{rank}A_{X}<n$. In particular, it is sufficient that each interior of a maximal cell of $\Delta$ contains a data point, possibly after adding minimal noise. From Theorem 3.5 together with Dostert and Jochemko (2023, Corollary 3.14) we obtain the following sufficient condition on the uniqueness of the maximum of $\mathcal{L}(\mathbf{a})$.
@@
Let $X_{0}$ and $X_{1}$ be sets of noisy data labeled with 00 and $1$ respectively, and such that for every maximal cell $\sigma\in\Delta$ there is at least one data point from $X_{0}$ and $X_{1}$ in the interior of $\sigma$. Then $\mathcal{L}(\mathbf{a})$ has a unique (possibly degenerate) maximum in the closed positive orthant $\mathbb{R}_{\geq 0}^{n}$.
@@
Let $X$ be a dataset and $\lambda_{0}>0$ be a rate parameter such that $\mathcal{L}(\lambda_{0},\mathbf{a})$ has a unique maximum $\mathbf{a}^{*}(\lambda_{0})$. Then $\mathcal{L}(\lambda,\mathbf{a})$ has a unique maximum for all $\lambda>0$, denoted $\mathbf{a}^{*}(\lambda)$, and the function $\lambda\mapsto\mathbf{a}^{*}(\lambda)$ is a straight line inside >$\mathbb{R}^{n}_{>0}$ approaching the origin.
[…]
@@
Given the monotonicity of the number of false positives and false negatives in the optimal solution, it is natural to ask if the $0/1$-loss, which is given by their sum, is convex in the sense that it first decreases and then increases with varying $\lambda$, without any “ups” and “downs”. However, this is in general not the case. A counterexample is given by Section A.2.
@@ 4 Geometry of the generalized parameter space
In the previous section, we have considered polyhedral fans whose cones have their apex at the origin, and varied the shapes of the stars defined on this fixed fan. In this section, we extend this framework by allowing translations. In Section 4.1 we first fix the shape of the star and only vary it by translation, whereas in Section 4.2 we investigate the space of both operations at the same time.
@@ 4.1 Translations of a fixed star
where $[\mathbf{x}-\mathbf{t}]^{\Delta}$ captures the nonzero coefficients $\mu_{i}$ such that $\mathbf{x}=\mathbf{t}+\mu_{i_{1}}\mathbf{v}_{i_{1}}+\ldots+\mu_{i_{k}}\mathbf{v}_{i_{k}}\in C+\mathbf{t}$, and $C+\mathbf{t}\in\Delta+\mathbf{t}$ is the unique cone of the translated fan containing $\mathbf{x}$. On the other hand, we have
@@
where $\mathbf{x}-C=\{\mathbf{x}-\mathbf{c}\mid\mathbf{c}\in C\}$ is the reflected cone $-C$ translated by the vector $\mathbf{x}$. We first analyze the behavior of $[\mathbf{x}-\mathbf{t}]^{\Delta}$ when varying $\mathbf{t}$.
[…]
@@
[…]
Given a fixed $\mathbf{a}\in\mathbb{R}_{>0}^{n}$, the data star arrangement of a dataset $X=\{(\mathbf{x}^{(i)},y^{(i)})\}_{i=1}^{m}$ is the arrangement of stars $-\operatorname{star}(\mathbf{a})+\mathbf{x}^{(i)}$ for $i=1,\dots,m$.
An example of the data star arrangement of a $2$-dimensional dataset, together with the level sets of the translational 0/1-loss is given in Section A.3 in Section A.3.
@@ 4.2 Translations and transformations of the star together
In Section 3 we have considered classification by a starshaped polyhedral set in which $\mathbf{t}\in\mathbb{R}^{d}$ is fixed (and assumed to be the origin), and varied $\mathbf{a}\in\mathbb{R}^{n}_{>0}$. In Section 4.1 we have fixed $\mathbf{a}\in\mathbb{R}^{n}_{>0}$ and varied $\mathbf{t}\in\mathbb{R}^{d}$. As a final step, we will now vary the tuple $(\mathbf{a},\mathbf{t})\in\mathbb{R}^{n}_{>0}\times\mathbb{R}^{d}$, i.e., both parameters simultaneously, and consider the classifiers
@@
We extend the 0/1-loss to be viewed as a function $\operatorname{err}(\mathbf{a},\mathbf{t})$ in variables $(\mathbf{a},\mathbf{t})\in\mathbb{R}^{n}_{>0}\times\mathbb{R}^{d}$. In contrast to Theorem 3.3 and Theorem 4.1, the (sub)level sets in this extended product of both parameter spaces are neither polyhedral nor do they have piecewise-linear boundary, but they are semialgebraic.
@@
The level sets and sublevel sets of the extended 0/1-loss on >$\mathbb{R}^{n}_{>0}\times\mathbb{R}^{d}$ are semialgebraic sets, i.e., finite unions and intersections of solutions to polynomial inequalities. The defining polynomials have degree at most $2$.
@@ 5 Conclusion
The presented framework has not been tested on real-world data yet. It remains an open question how to optimally select the parameters, such the underlying fan, the translation vector and the rate parameter, in a manner tailored to the specific problem at hand.
6 Acknowledgements
We would like to thank Maria Dostert and Mariel Supina for many fruitful discussions. We also want to thank Roland Púček for insightful conversations. MCB and KJ were partially supported by the Wallenberg AI, Autonomous Systems and Software Program funded by the Knut and Alice Wallenberg Foundation. KJ was furthermore supported by grant nr 2018-03968 of the Swedish Research Council as well as the Göran Gustafsson Foundation. MCB was partially supported by the SPP 2458 "Combinatorial Synergies", funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 521249412.
@@ Appendix A Examples
A.1 Examples of polyhedral fans (Section 2)
[…]
$[\mathbf{x}]^{\Diamond}\ =\ \max\left(\mathbf{0},\mathbf{x}^{T}A_{\diamond}\right)\,$
where the maximum is taken coordinatewise. The $2$-dimensional fan $\diamond$ and the function $[\mathbf{x}]^{\diamond}$ restricted to each maximal cone is depicted in Figure 3(a).

### v2

@@ 3.1 Geometry of the 0/1-loss
We seek to understand the geometry inside the parameter space $\mathbb{R}_{>0}^{n}=\{\mathbf{a}:\mathbf{a}>0\}$. For an example which illustrates all definitions and results stated in this and the following subsection (Sections˜3.1 and 3.2), we refer to Section˜A.2 in Section˜A.2.
@@
A direct consequence of Section˜3.1 is that the false positives, the false negatives and thus also the discrete loss is constant on the half-open chambers. In general, since every chamber is convex and corresponds to a unique classification, the set of all parameters $\mathbf{a}$ that perfectly separate the data points is convex.
Corollary 3.3.
@@
Section˜3.1 implies that if the dataset is separable, i.e., if $L(\operatorname{err},0)\neq\emptyset$, then the set of perfect classifiers $L(\operatorname{err},0)$ is a convex set. Under the same assumption, we can make a similar statement as Theorem˜3.4 about the sublevel sets of the discrete loss function.
@@
Theorem˜3.5 shows that if the data is separable, then the sublevel sets of $\operatorname{err}$ are star-convex, but not necessarily convex in the usual sense (see also Section˜A.2 in Section˜A.2). Much stronger, in the case of non-separable data, sublevel sets and even the set of minimizers of the discrete loss can be disconnected. One example for this scenario is the point configuration from Section˜A.2, with labeling $0,0,1,1,1,1,0,0$. Here, the minimum of the discrete loss-function is $4$, attained in four non-neighboring half-open chambers.
To summarize the results in this subsection, our geometric analysis reveals that sublevel sets under the 0/1-loss decompose into convex chambers within a hyperplane arrangement (Section˜3.1), while simultaneously exhibiting star‑convexity with respect to any optimal point in the chamber which minimizes the loss (Theorems˜3.4 and 3.5). These theorems establish that while global optimization over the 0-1 loss is hard due to the combinatorial complexity of the overall non‑convex landscape, the local landscape is well-connected. In particular, for separable data these results show that local optimization methods can between neighboring chambers along straight lines, improving in each step without getting stuck at local minima. For non-separable data, a similar behavior is still exhibited by false positives and false negatives.
3.2 Log-likelihood loss
In Section˜3.1 we have observed that the 0/1-loss or discrete loss admits discrete geometric structures in the parameter space which are governed by an affine hyperplane arrangement. However, the discrete loss function is difficult to compute in practice. We thus propose an alternative loss-function for practical purposes. The log-likelihood loss function turns out to be well-suited for optimization procedures due to concavity. The definitions and results stated in this subsection are exemplified in Section˜A.2 in Section˜A.2.
@@ 3.2 Log-likelihood loss
The computation of the maximum likelihood estimator, the maximizer of the log-likelihood loss-function, can be summarized by the following algorithm.
@@
Convex (and thus also concave) functions can be optimized in polynomial time. Moreover, the description in Section˜2 implies that $A_{X}$ can be determined in polynomial time. This implies the following.
Corollary 3.7.
Algorithm˜1 can be computed in polynomial time in the size of the input data.
Strictly speaking, for Algorithm˜1 to be a convex program, we need to consider the closed positive orthant. Any solution on the boundary corresponds to a degenerate star, where star-defining points on rays move to infinity. This degenerate case will be treated in Theorem˜3.9.
Similarly to Section˜3.1, we now consider the superlevel sets of the log-likelihood loss. For given $t\in\mathbb{R}$, the superlevel $S_{\mathcal{L}}(t)$ of $\mathcal{L}$ is defined as
@@
We now analyze cases in which the maximum of the log-likelihood loss is unique. For this, consider the positively labeled subdataset $X_{1}=\{\mathbf{x}^{(i)}\colon y^{(i)}=1\}\subset X$ and let $A_{X_{1}}$ be the submatrix of $A_{X}$ composed of all rows $\mathbf{x}^{(i)}]$ corresponding to $\mathbf{x}^{(i)}\in X_{1}$; similarly we define $A_{X_{0}}$ for $X_{0}=X\setminus X_{1}$. Then we have the following sufficient condition for a unique maximum of $\mathcal{L}(\mathbf{a})$.
@@
The log-likelihood loss function $\mathcal{L}(\mathbf{a})$ is strictly concave if the matrix $A_{X_{1}}$ has rank $n$. Furthermore, if also the rank of $A_{X_{0}}$ equals $n$, then $\mathcal{L}(\mathbf{a})$ has a unique (possibly degenerate) maximum in the closed positive orthant $\mathbb{R}_{\geq 0}^{n}$.
Matrices of the form $A_{X}$ also appear in the study of reconstruction of polytopes with fixed facet directions from support function evaluations. LearningPolytopes showed that such a reconstruction of a polytope is unique if and only if $\operatorname{rank}A_{X}=n$. In other words, $A_{X}\in(\mathbb{R}^{n\times m})\setminus\mathbb{V}$ where $\mathbb{V}$ is the algebraic variety encoding that $\operatorname{rank}A_{X}<n$. In particular, it is sufficient that each interior of a maximal cell of $\Delta$ contains a data point, possibly after adding minimal noise. From Theorem˜3.9 together with LearningPolytopes we obtain the following sufficient condition on the uniqueness of the maximum of $\mathcal{L}(\mathbf{a})$.
@@
Let $X_{0}$ and $X_{1}$ be sets of noisy data labeled with 0 and $1$ respectively, and such that for every maximal cell $\sigma\in\Delta$ there is at least one data point from $X_{0}$ and $X_{1}$ in the interior of $\sigma$. Then $\mathcal{L}(\mathbf{a})$ has a unique (possibly degenerate) maximum in the closed positive orthant $\mathbb{R}_{\geq 0}^{n}$.
@@
Let $X$ be a dataset and $\lambda_{0}>0$ be a rate parameter such that $\mathcal{L}(\lambda_{0},\mathbf{a})$ has a unique maximum $\mathbf{a}^{*}(\lambda_{0})$. Then $\mathcal{L}(\lambda,\mathbf{a})$ has a unique maximum for all $\lambda>0$, denoted $\mathbf{a}^{*}(\lambda)$, and the function $\lambda\mapsto\mathbf{a}^{*}(\lambda)$ is a straight line inside $\mathbb{R}^{n}_{>0}$ approaching the origin.
Corollary 3.12.
@@
Given the monotonicity of the number of false positives and false negatives in the optimal solution, it is natural to ask if the $0/1$-loss, which is given by their sum, is convex in the sense that it first decreases and then increases with varying $\lambda$, without any “ups” and “downs”. However, this is in general not the case. A counterexample is given by Section˜A.2.
@@ 4 Geometry of the generalized parameter space
In the previous section, we have considered polyhedral fans whose cones have their apex at the origin, and varied the shapes of the stars defined on this fixed fan. In this section, we extend this framework by allowing translations. In Section˜4.1 we first fix the shape of the star and only vary it by translation, whereas in Section˜4.2 we investigate the space of both operations at the same time.
@@ 4.1 Translations of a fixed star
where $\mathbf{x}-\mathbf{t}]^{\Delta}$ captures the nonzero coefficients $\mu_{i}$ such that $\mathbf{x}=\mathbf{t}+\mu_{i_{1}}\mathbf{v}_{i_{1}}+\ldots+\mu_{i_{k}}\mathbf{v}_{i_{k}}\in C+\mathbf{t}$, and $C+\mathbf{t}\in\Delta+\mathbf{t}$ is the unique cone of the translated fan containing $\mathbf{x}$. On the other hand, we have
@@
where $\mathbf{x}-C=\{\mathbf{x}-\mathbf{c}\mid\mathbf{c}\in C\}$ is the reflected cone C-C translated by the vector $\mathbf{x}$. We first analyze the behavior of $\mathbf{x}-\mathbf{t}]^{\Delta}$ when varying $\mathbf{t}$.
Proposition 4.1.
@@
Definition 4.2.
Given a fixed $\mathbf{a}\in\mathbb{R}_{>0}^{n}$, the data star arrangement of a dataset $X=\{(\mathbf{x}^{(i)},y^{(i)})\}_{i=1}^{m}$ is the arrangement of stars $\operatorname{star}(\mathbf{a})+\mathbf{x}^{(i)}$ for $i=1,\dots,m$.
An example of the data star arrangement of a $2$-dimensional dataset, together with the level sets of the translational 0/1-loss is given in Section˜A.3 in Section˜A.3.
@@ 4.2 Translations and transformations of the star together
In Section˜3 we have considered classification by a starshaped polyhedral set in which $\mathbf{t}\in\mathbb{R}^{d}$ is fixed (and assumed to be the origin), and varied $\mathbf{a}\in\mathbb{R}^{n}_{>0}$. In Section˜4.1 we have fixed $\mathbf{a}\in\mathbb{R}^{n}_{>0}$ and varied $\mathbf{t}\in\mathbb{R}^{d}$. As a final step, we will now vary the tuple $(\mathbf{a},\mathbf{t})\in\mathbb{R}^{n}_{>0}\times\mathbb{R}^{d}$, i.e., both parameters simultaneously, and consider the classifiers
@@
We extend the 0/1-loss to be viewed as a function $\operatorname{err}(\mathbf{a},\mathbf{t})$ in variables $(\mathbf{a},\mathbf{t})\in\mathbb{R}^{n}_{>0}\times\mathbb{R}^{d}$. In contrast to Theorem˜3.5 and Theorem˜4.3, the (sub)level sets in this extended product of both parameter spaces are neither polyhedral nor do they have piecewise-linear boundary, but they are semialgebraic.
@@
The level sets and sublevel sets of the extended 0/1-loss on $\mathbb{R}^{n}_{>0}\times\mathbb{R}^{d}$ are semialgebraic sets, i.e., finite unions and intersections of solutions to polynomial inequalities. The defining polynomials have degree at most $2$.
@@
5 Experiments
We conducted small-scale experiments where we tested Algorithm 1, implemented in SageMath 10.5 (sagemath), on two-dimensional synthetic data. The computations were done on a MacBook Pro equipped with an M2 Pro chip and $32$ GB of RAM. For comparison, we also applied several standard binary classification methods leading to convex optimization problems on the same dataset, as well as a ReLU neural network. The computation running time ranged from few seconds to one hour.
5.1 Data
Figure 3(a) illustrates $500$ data points sampled from a given star-shaped region (in green) defined on eight rays. The data was generated as follows: we randomly selected the $x$- and $y$-coordinates of all points from the interval [1,1][-1,1] using a uniform distribution and discarded any resulting points $(x,y)$ lying outside the unit circle. This was done to achieve a near rotational symmetry of the data set. For each remaining point, we then checked whether it lies inside or outside the star-shaped region. The corresponding label was assigned accordingly, with a 90% probability of being correct.
5.2 Results
In the first experiment, we used Algorithm 1 together with the same eight-ray fan structure to predict the labels of the synthetic data described above. We compared binary classification of our algorithm with multiple standard classification models. In the following we give a description of our results. Illustrations can be found in Appendix B
Running Algorithm˜1 on the synthetic data set, the optimal value of the regularization parameter was found to be approximately $\lambda=0.83$, yielding an accuracy of $0.852$. The resulting optimal star classifier is shown in Figure 10(a). For comparison, we also tested standard implementations of SVMs (with linear, polynomial, RBF, and sigmoid kernels), logistic regression, and a ReLU neural network with two hidden layers of sizes $5$ and $2$, respectively. The SVMs with linear and polynomial kernels, as well as logistic regression, performed poorly, assigning all points to the same class, thereby achieving an accuracy of $0.718$. The SVM with a sigmoid kernel performed even worse. In contrast, the SVM with an RBF kernel and the neural network achieved better results, with accuracies of $0.78$ and $0.824$, respectively. See Figure 3(b) for a summary of the results.
(a) Synthetic data.
Model Accuracy
SVM sigmoid 0.534
SVM linear 0.718
SVM poly 0.718
Logistic reg 0.718
SVM RBF 0.78
Neural net 0.824
Algorithm˜1 0.852
(b) Results.
Figure 3: Synthetic data and accuracy of classification for tested models.
In a further experiment, we ran Algorithm 1 on the same dataset but with different underlying fans as input. Specifically, we considered both a refinement and a coarsening of the original fan with eight rays, as depicted in Figure˜11. In the case of the refined fan, the decision boundary remained almost unchanged and the accuracy improved marginally. For the coarsened fan, the shape of the decision boundary changed considerably and the accuracy became significantly worse depending on which ray was removed; see Figure˜12 for an illustration if the starshaped sets. These results support the following heuristic for fan selection in two dimensions: start with a small number of rays and iteratively refine the fan by adding more rays. If the additional rays do not produce significant new dents in the boundary, they can be safely discarded.
@@ 6 Conclusion
The presented framework has been tested only on very few example data sets in two dimensions. It remains an open question how to optimally select the parameters, such the underlying fan, the translation vector and the rate parameter, in a manner tailored to the specific problem at hand.
Acknowledgements
We would like to thank Maria Dostert and Mariel Supina for many fruitful discussions. We also want to thank Roland Púček for insightful conversations. MB and KJ were supported by the Wallenberg AI, Autonomous Systems and Software Program funded by the Knut and Alice Wallenberg Foundation. KJ was furthermore supported by grant nr 2018-03968 and nr 2023-04063 of the Swedish Research Council as well as the Göran Gustafsson Foundation. MB was furthermore supported by the SPP 2458 "Combinatorial Synergies", funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation).
@@ Appendix A Examples
A.1 Examples of polyhedral fans (Section˜2)
Example A.1 (Kites).
For $1\leq i\leq d$, let $\mathbf{e}_{i}\in\mathbb{R}^{d}$ be the vector with $1$ in coordinate $i$ and 0 in all other coordinates. The coordinate hyperplanes $\{\mathbf{x}\in\mathbb{R}^{d}:\langle\mathbf{e}_{i},\mathbf{x}\rangle=0\}$ divide $\mathbb{R}^{d}$ into chambers, and these chambers along with all of their faces form a simplicial fan $\Diamond$ with generators $\{\pm\mathbf{e}_{i}\}_{1\leq i\leq d}$. Every star arising from this fan is necessarily convex, and such stars are called kites. Consider the $(d\times 2d)$-matrix $A_{\diamond}=\left(\mathbf{e}_{1}\,,-\mathbf{e}_{1}\,,\mathbf{e}_{2}\,,-\mathbf{e}_{2}\,,\dots\,,\mathbf{e}_{d}\,,-\mathbf{e}_{d}\right)$. Given $\mathbf{x}\in\mathbb{R}^{d}$, the vector $\mathbf{x}]^{\Diamond}$ is obtained as
$\mathbf{x}]^{\Diamond}\ =\ \max\left(\mathbf{0},\mathbf{x}^{T}A_{\diamond}\right)\,$
where the maximum is taken coordinatewise. The $2$-dimensional fan $\diamond$ and the function $\mathbf{x}]^{\diamond}$ restricted to each maximal cone is depicted in Figure˜4(a).